OpenStudy (anonymous):

First find the slope and then an equation of the tangent line to the graph of √x at x = 81.

OpenStudy (anonymous):

find the slope by taking the derivative and replacing x by 81

OpenStudy (anonymous):

Would the derivative be (1/2)x^(-1/2)?

OpenStudy (anonymous):

you get \[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\] so your slope is \[\frac{1}{18}\]

OpenStudy (anonymous):

What's "d"?

OpenStudy (anonymous):

yes it is true that you can write \[\frac{1}{2}x^{-\frac{1}{2}}\] but that form is useless for computing

OpenStudy (anonymous):

that just means the derivative. write \[y=\sqrt{x}\] \[y'=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

How do I write the equation?

OpenStudy (anonymous):

use the "point-slope" formula with \[m=\frac{1}{18},x_1=81,y_1=9\]

OpenStudy (anonymous):

Thanks